# Benchmarking Large Vision-Language Models via Directed Scene Graph for Comprehensive Image Captioning

### v1

@@ 4.2 Comprehensive Image Caption Evaluation
4.3 Fine-grained Object VQA
We show the proportion of pixels of objects omitted in the generated captions in Fig. 7. Most of these ignored objects are tiny, the segmentation map of which only occupies less than 5% of the global image. To analyze the correlation between LVLMs’ perception of fine-grained objects and their qualities of generating detailed captions, we evaluate the 9 LVLMs on the fine-grained object VQA including CompreQA-P and CompreQA-Cap. The evaluation for GPT-4V is lacking because its API has been deprecated after annotating the final versions of this VQA. The results in Tab. 3 show that InternVL-Chat-V1-5 [9] achieves the best performance among LVLMs benefiting from a powerful visual encoder with 6B parameters. LLaVA-Next-34B [27] and GPT-4o [32] rank second and third, respectively. In CompreQA-P, the hallucination created by non-existent objects results in InstructBLIP-7B [11] and MiniGPT4-v2 [49] only achieving the accuracy rates comparable to, or even worse than random guessing (i.e., 50%). The human performance also surpasses all LVLMs on both CompreQA-P and CompreQA-Cap, with an accuracy over 96%, underscoring the rationality of the fine-grained object VQA task in CompreCap. In general, the ranking of LVLMs on this VQA task aligns with their ranking in terms of comprehensive caption quality.
@@ 5 Conclusion
Appendix A More Experimental Analysis
A.1 Analysis on the effect of components of LVLMs
@@
A.2 Consistency with human evaluation scores
@@
InstructBLIP-7B [11] 69.93 56.20 1.89±$\pm\emph{0.00}$ 2.53±$\pm\emph{0.00}$ 42.03±$\pm\emph{0.14}$ 2.36
MiniGPT4-v2 [49] 350.42 56.74 1.86±$\pm\emph{0.00}$ 1.88±$\pm\emph{0.01}$ 43.03±$\pm\emph{0.19}$ 1.83
LLaVA-1.5-13B [28] 86.97 59.86 2.01±$\pm\emph{0.01}$ 2.59±$\pm\emph{0.00}$ 43.81±$\pm\emph{0.25}$ 2.39
ShareGPT4V-13B [8] 155.91 67.88 2.40±$\pm\emph{0.01}$ 2.69±$\pm\emph{0.00}$ 55.86±$\pm\emph{0.17}$ 3.28
LLaVA-Next-llama3-8B [27] 168.99 70.22 2.48±$\pm\emph{0.00}$ 2.72±$\pm\emph{0.01}$ 56.95±$\pm\emph{0.08}$ 3.34
miniGemini-HD-34B [22] 173.71 70.70 2.48±$\pm\emph{0.00}$ 2.70±$\pm\emph{0.00}$ 57.20±$\pm\emph{0.11}$ 3.37
InternVL-Chat-V1-5 [9] 115.22 70.56 2.50±$\pm\emph{0.00}$ 2.87±$\pm\emph{0.00}$ 57.58±$\pm\emph{0.09}$ 3.42
LLaVA-Next-34B [27] 179.24 72.86 2.59±$\pm\emph{0.00}$ 2.79±$\pm\emph{0.00}$ 58.49±$\pm\emph{0.15}$ 3.64
GPT-4V [44] 202.06 72.31 2.52±$\pm\emph{0.00}$ 2.73±$\pm\emph{0.00}$ 57.27±$\pm\emph{0.14}$ 3.63
GPT-4o [32] 108.20 72.78 2.58±$\pm\emph{0.00}$ 2.93±$\pm\emph{0.00}$ 57.54±$\pm\emph{0.23}$ 3.68
Human 133.61 77.62 2.78±$\pm\emph{0.00}$ 2.99±$\pm\emph{0.00}$ 59.58±$\pm\emph{0.16}$ 4.0
@@ A.3 Compare with directly scoring with Llama3
A.4 Analysis on the influence of LLM evaluator
A.5 Analysis on the effect of parsers
@@ A.10 Analysis on the failure cases in the fine-grained object VQA
Appendix B Data Comparison with MSCOCO
@@
Appendix C Data Samples from CompreCap dataset
Appendix D Discussion about More Long Caption Datasets
@@
Appendix E Limitations and Broader Impact

### v3

@@ 5 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants 62225207, 62436008, 62306295, 62476260 and the Fundamental Research Funds for the Central Universities under Grant WK2100000057.
@@ Acknowledgments
Appendix A More Experimental Analysis
@@
Evaluation
InstructBLIP-7B [12] 69.93 56.20 1.89±$\pm\emph{0.00}$ 2.53±$\pm\emph{0.00}$ 42.03±$\pm\emph{0.14}$ 48.16 2.36
MiniGPT4-v2 [55] 350.42 56.74 1.86±$\pm\emph{0.00}$ 1.88±$\pm\emph{0.01}$ 43.03±$\pm\emph{0.19}$ 42.28 1.83
LLaVA-1.5-13B [31] 86.97 59.86 2.01±$\pm\emph{0.01}$ 2.59±$\pm\emph{0.00}$ 43.81±$\pm\emph{0.25}$ 50.32 2.39
ShareGPT4V-13B [9] 155.91 67.88 2.40±$\pm\emph{0.01}$ 2.69±$\pm\emph{0.00}$ 55.86±$\pm\emph{0.17}$ 55.56 3.28
LLaVA-Next-llama3-8B [30] 168.99 70.22 2.48±$\pm\emph{0.00}$ 2.72±$\pm\emph{0.01}$ 56.95±$\pm\emph{0.08}$ 56.91 3.34
miniGemini-HD-34B [25] 173.71 70.70 2.48±$\pm\emph{0.00}$ 2.70±$\pm\emph{0.00}$ 57.20±$\pm\emph{0.11}$ 56.88 3.37
InternVL-Chat-V1-5 [10] 115.22 70.56 2.50±$\pm\emph{0.00}$ 2.87±$\pm\emph{0.00}$ 57.58±$\pm\emph{0.09}$ 58.48 3.42
LLaVA-Next-34B [30] 179.24 72.86 2.59±$\pm\emph{0.00}$ 2.79±$\pm\emph{0.00}$ 58.49±$\pm\emph{0.15}$ 58.85 3.64
GPT-4V [47] 202.06 72.31 2.52±$\pm\emph{0.00}$ 2.73±$\pm\emph{0.00}$ 57.27±$\pm\emph{0.14}$ 57.74 3.63
GPT-4o [35] 108.20 72.78 2.58±$\pm\emph{0.00}$ 2.93±$\pm\emph{0.00}$ 57.54±$\pm\emph{0.23}$ 60.05 3.68
Human 133.61 77.62 2.78±$\pm\emph{0.00}$ 2.99±$\pm\emph{0.00}$ 59.58±$\pm\emph{0.16}$ 62.99 4.0
A.1 Analysis on the effect of components of LVLMs
A.2 Consistency with human evaluation scores
@@ A.3 Compare with directly scoring with Llama3
We select the human performance, two top-performing models in Tab. 2 of the manuscript paper, GPT-4o and LLaVA-Next-34B, as well as a medium-performing model, ShareGPT4V13B, and a low-performing model, InstructBLIP7B, for comparative analysis. With the human assessment scores of the image captions generated by both LVLMs and human, the metrics derived from our method faithfully align with human evaluation scores compared to those directly obtained from Llama3, as shown in Fig. S1. The high consistency with human judgment emphasizes the credibility and practicality of our CompreCap, proving its promising prospects in LVLMs evaluation.
A.4 Ability of Llama3 to Distinguish Caption Quality
A.5 Analysis on the influence of LLM evaluator
@@
Figure S1: Our method demonstrates a high consistency with human evaluation scores across all models, whereas the results directly produced by Llama3 exhibit conflicts. This indicates that the annotations of the directed scene graph facilitate more accurate matching at the levels of objects, attributes and relations during comprehensive caption assessment.
Figure S5: The human performance exceed all LVLMs on the scores of objects, attributes and relations with our evaluation method. Compared to directly scoring detailed captions with Llama3, the annotations of the directed scene graph provide more precise references across various dimensions.
A.6 Analysis on the effect of parsers
@@ A.10 Analysis on the failure cases in the fine-grained object VQA
Table S8: Comparison with similar benchmarks. Visual-Genome includes VG-Attributes and VG-Relations.
Length Mask Object Attribute Relation
COCO-Stuff 10 - -
COCO-OOD 10 - -
SugarCrepe 11 - - -
Visual-Genome 92 -
CompreCap 172
Appendix B Comparison with Similar Benchmarks
Here we compare several related benchmarks in Tab. S8, especially Visual-Genome (VG) [21] and other COCO-based benchmarks. COCO-Stuff/OOD [19, 7] have object and mask annotations, while SugarCrepe generates various forms of hard negatives based on COCO’s short captions (averaged length is about 10 words). VG has object, attribute, and relation annotation with a bounding box, which has about 92 lengthy captions per image. Our CompreCap has a more detailed caption (about 172) and more dense mask annotation than VG, where the bounding box coverage has about 99.23% vs. 85.26% of VG. For comprehensive image captioning, a dataset with wider coverage and a more detailed description is important. Moreover, based on our dataset, we also propose an evaluation metric for comprehensive image captioning.
They all lack annotations of object-bound attributes and directional relations. VG is the closest to our dataset, but its relatively brief annotations limit the diversity of attribute and relation descriptions, and its lower coverage cannot ensure the inclusion of all main objects in images, hindering the comprehensive evaluation of detailed captions. Moreover, VG does not propose a method to assess coherent long image captions. Instead, it makes the model output captions for different image regions sequentially and still uses traditional n-gram evaluation metrics.
The characteristics of CompreCap are highlighted as follows: 1. Comprehensive annotation: CompreCap has object, attribute, and relation annotations with pixel-level masks; 2. More detailed caption: CompreCap has more detailed human-annotated captions, with an average of 172 words vs. VG’s 92; 3. Wider coverage: The pixel coverage of CompreCap is 95.83%. Its bounding box coverage is 99.23% vs. 85.26% of VG; 4. Evaluation method for long caption of image: Compared to other benchmarks, CompreCap proposes a method to comprehensively evaluate the visual context in detailed captions.
Appendix C Data Comparison with MSCOCO
@@
Appendix D Data Samples from CompreCap dataset
Appendix E Discussion about More Long Caption Datasets
Appendix F Information of Human Participants
In the construction and experimentation process of CompreCap, three stages required the involvement of human experts: data annotation, manual captioning of images for evaluating human performance, and manual scoring of captions generated by models and humans. We engaged a total of 50 participants from universities and a crowdsourcing platform, dividing them into groups of 20, 10, and 20 participants to respectively take part in these three stages, ensuring no overlap of personnel across the stages. These participants are aged 22 to 45, with backgrounds in linguistics, computer science, and mathematics.
During the manual captioning of images, the 10 human experts each described 56 different images as comprehensively as possible. They were not specifically instructed to focus on objects, attributes, or relations in their descriptions. During the manual evaluation scoring stage, the 20 experts independently scored all 6,160 captions generated by both models and humans according to their personal assessment criteria. The average score was used as the final score for both the models and human performance.
Appendix G Limitations and Broader Impact